# RoFt-Mol: Benchmarking Robust Fine-Tuning with Molecular Graph Foundation Models

### v2

@@ Appendix E Dataset Statistics
Dataset Evaluation Metrics Task # Tasks # Molecules BBBP [Martins et al., 2012] AUC Classification 1 2,039 Tox21 AUC Classification 12 7,831 ToxCast [Richard et al., 2016] AUC Classification 617 8,576 Sider [Kuhn et al., 2016] AUC Classification 27 1,427 ClinTox [Gayvert et al., 2016] AUC Classification 2 1,478 MUV [Rohrer and Baumann, 2009] AUC Classification 17 93,087 HIV Zaharevitz [2015] AUC Classification 1 41,127 Bace [Subramanian et al., 2016] AUC Classification 1 1,513 Esol [Delaney, 2004] RMSE Regression 1 1,128 Lipo [Gaulton et al., 2012] RMSE Regression 1 4,200 Malaria [Gamo et al., 2010] RMSE Regression 1 9,999 CEP [Hachmann et al., 2011] RMSE Regression 1 29,978
@@ Appendix F Details of Experimental Implementation
Fine-tuning Implementations. We keep the same training configurations across all the downstream datasets, pre-training models, and fine-tuning strategies, following Hu et al. [2020a]. Specifically, for each distinct setting, we fine-tune the pre-training models with 5 random seeds (0-4). We use a batch size of 32 and a dropout rate of 0.5. For each dataset, We train models for 100 epochs and report the test performance when the optimal validation performance is achieved.
@@ G.1 Comparisons over pre-trained models
Lastly, note that in the few-shot setting, GraphSTM underperforms other evaluated models in self-supervised pretraining. This is mainly because GraphSTM’s GNN encoder was specifically pretrained with graph-text alignment to enhance multi-modal tasks like structure-text retrieval. Therefore, the encoder would retain features optimized for cross-modal alignment rather than purely graph structural information. Since the downstream tasks in our benchmark do not involve text, the randomly initialized task head struggles to effectively utilize these features with limited data, whereas other models provide a more direct and task-relevant representation, leading to better performance in low-data scenarios.
@@ G.2 Comparisons over traditional method
G.3 Additional findings
@@
For few-shot fine-tuning with 50 and 100 samples (c.f., Fig. 2b and 2d), LP surpasses full FT in random and scaffold splits, differing from self-supervised pre-training discussed in (1a). This again supports the claim that directly adopting molecular representations from supervised pre-training retain useful knowledge for downstream tasks. But interestingly, this does not hold for size splits. We believe it is due to the susceptibility of graph level tasks under size shift, as noted in prior OOD studies [Zou et al., 2024]. Namely, the prediction head tends to overfit to the mapping from representations to output labels with molecules in a specific range of sizes, and thus cannot generalize to OOD molecules of different sizes.
@@ Appendix H Additional Experimental Results
In this section, we present complementary baseline results that are not shown in the main text due to space limit. Table 1 is a summary of all pre-trained models we test on and their corresponding result tables for reference.
The complete table including all few-shot fine-tuning results for DWiSE-FT are in Table LABEL:table:new_method_appendix.
@@
Split Methods ClinTox BBBP Bace HIV MUV Sider Tox21 ToxCast Avg Avg-F Avg-R Self-supervised Pre-training (MoleculeSTM) Random Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS Scaffold Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS Size Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS Supervised Pre-training (Graphium-Large) Random Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS Size Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS
Split Methods Self-supervised Pre-training (MoleculeSTM) Supervised Pre-training (Graphium-Large) Esol Lipo Malaria Cep Avg-R Avg-R∗ Esol Lipo Malaria Cep Avg-R∗ Random Full-FT 2 4 LP 8 8 Surgical-FT 6 1 LP-FT 3 2 WiSE-FT 5 3 -SP 4 7 Feature-map 7 5 BSS 1 6 Scaffold Full-FT 5 5 LP 8 7 Surgical-FT 6 3 LP-FT 4 4 WiSE-FT 2 6 -SP 3 8 Feature-map 7 2 BSS 1 1 Size Full-FT 2 2 LP 8 8 Surgical-FT 6 5 LP-FT 3 4 WiSE-FT 5 1 -SP 4 6 Feature-map 7 7 BSS 1 3
@@
Split Methods Self-supervised Pre-training (MoleculeSTM) Supervised Pre-training (Graphium-Large) Esol Lipo Malaria CEP Avg-R Avg-R∗ Esol Lipo Malaria CEP Avg-R∗ Fewshot-50 Random Full-FT 6 7 LP 8 8 Surgical-FT 7 4 LP-FT 4 1 WiSE-FT 5 3 -SP 1 5 Feature-map 3 2 BSS 2 6 Scaffold Full-FT 6 7 LP 8 1 Surgical-FT 7 2 LP-FT 4 8 WiSE-FT 5 4 -SP 2 6 Feature-map 3 5 BSS 1 3 Size Full-FT 7 4 LP 8 8 Surgical-FT 6 1 LP-FT 4 7 WiSE-FT 5 5 -SP 1 6 Feature-map 3 3 BSS 2 2 Fewshot-100 Random Full-FT 6 6 LP 8 4 Surgical-FT 4 2 LP-FT 7 7 WiSE-FT 5 3 -SP 1 5 Feature-map 2 1 BSS 3 8 Scaffold Full-FT 6 7 LP 8 3 Surgical-FT 5 1 LP-FT 7 8 WiSE-FT 4 5 -SP 1 6 Feature-map 3 2 BSS 2 4 Size Full-FT 7 6 LP 8 8 Surgical-FT 6 1 LP-FT 5 7 WiSE-FT 4 5 -SP 1 3 Feature-map 2 4 BSS 3 2 Fewshot-500 Random Full-FT 6 5 LP 8 7 Surgical-FT 7 1 LP-FT 5 2 WiSE-FT 4 3 -SP 1 8 Feature-map 3 4 BSS 2 6 Scaffold Full-FT 6 6 LP 8 3 Surgical-FT 5 2 LP-FT 7 4 WiSE-FT 4 5 -SP 1 7 Feature-map 3 1 BSS 2 8 Size Full-FT 7 7 LP 8 8 Surgical-FT 4 4 LP-FT 5 3 WiSE-FT 6 5 -SP 1 1 Feature-map 2 6 BSS 3 2

### v3

@@ Appendix E Dataset Statistics
Dataset Evaluation Metrics Task # Tasks # Molecules BBBP [martins2012bayesian] AUC Classification 1 2,039 Tox21 AUC Classification 12 7,831 ToxCast [richard2016toxcast] AUC Classification 617 8,576 Sider [kuhn2016sider] AUC Classification 27 1,427 ClinTox [gayvert2016data] AUC Classification 2 1,478 MUV [rohrer2009maximum] AUC Classification 17 93,087 HIV zaharevitz2015aids AUC Classification 1 41,127 Bace [subramanian2016computational] AUC Classification 1 1,513 Esol [delaney2004esol] RMSE Regression 1 1,128 Lipo [gaulton2012chembl] RMSE Regression 1 4,200 Malaria [gamo2010thousands] RMSE Regression 1 9,999 CEP [hachmann2011harvard] RMSE Regression 1 29,978
@@ Appendix F Details of Experimental Implementation
Fine-tuning Implementations. We keep the same training configurations across all the downstream datasets, pre-training models, and fine-tuning strategies, following hu2020strategies. Specifically, for each distinct setting, we fine-tune the pre-training models with 5 random seeds (0-4). We use a batch size of 32 and a dropout rate of 0.5. For each dataset, We train models for 100 epochs and report the test performance when the optimal validation performance is achieved.
@@ G.1 Comparisons over pre-trained models
The conclusions presented in Section 4 generalize well to models pre-trained on large-scale datasets, such as GraphGPS (pre-trained on PCQM4MV2) and Graphium-Large (trained on the LargeMix dataset containing hundreds of millions of labeled molecular graphs). In Section 4.3 and Appendix G.1, we analyze the consistency of trends across all six pre-trained models in our benchmark. Below, we summarize key observations that hold true for models trained on large-scale data:
Supervised pre-training on large datasets leads to stronger downstream performance, particularly in few-shot settings. This aligns with our main conclusion in Section 4.2 (Q2), where we compare supervised and self-supervised pre-training. Models like GraphGPS and Graphium-Large consistently outperform self-supervised models such as Mole-BERT and GraphMAE under the same fine-tuning protocols.
Representation-based fine-tuning methods (e.g., Feature-map and BSS) remain top-performing strategies for supervised pre-trained models on large-scale datasets, consistent with Finding 3. This trend holds across both classification and regression tasks, and across different dataset splits.
Partial fine-tuning methods (e.g., LP and Surgical-FT) continue to underperform in few-shot settings. This observation supports Finding 2, and reflects their tendency to underfit in data-scarce regimes, even when the underlying pre-trained model is strong.
@@ G.2 Comparisons over traditional method
G.3 Study on parameter efficient fine-tuning methods
As an additional study over parameter efficient fine-tuning method, we incorporate the LoRA [hu2022lora] results for GraphGPS under the scaffold split across three regimes: Fewshot-50, Fewshot-500, and non-fewshot. The results are shown in the table LABEL:table:lora.
Across both classification and regression tasks, LoRA falls short of full fine-tuning in roughly two-thirds of cases, with the gap widening for regression when more samples are available. This pattern is unsurprising since more challenging tasks and larger downstream datasets generally require updating a greater number of parameters. In the instances where LoRA does outperform full-FT, its results typically lie between standard full fine-tuning and the strongest fine-tuning baselines. Notably, under the Fewshot-50 regression setting, LoRA occasionally matches or even exceeds the best benchmarked fine-tuning methods, highlighting its potential in extremely low-data scenarios.
G.4 Additional study of DWiSE-FT on other pretrained models
We additionally test DWiSE-FT on other pretrained model like GraphGPS. As shown in Table LABEL:table:dwise-gps, we report the results of fewshot fine-tuning with 50 samples under scaffold and size splits. These results show that DWiSE-FT not only significantly improves over WiSE-FT and -SP, but also matches or exceeds the best-performing method (TOP) in some cases. This demonstrates that DWiSE-FT remains effective even under supervised pre-training, including on models like GraphGPS where traditional weight-based methods struggle.
G.5 Additional findings
@@
For few-shot fine-tuning with 50 and 100 samples (c.f., Fig. 2b and 2d), LP surpasses full FT in random and scaffold splits, differing from self-supervised pre-training discussed in (1a). This again supports the claim that directly adopting molecular representations from supervised pre-training retain useful knowledge for downstream tasks. But interestingly, this does not hold for size splits. We believe it is due to the susceptibility of graph level tasks under size shift, as noted in prior OOD studies [zougess]. Namely, the prediction head tends to overfit to the mapping from representations to output labels with molecules in a specific range of sizes, and thus cannot generalize to OOD molecules of different sizes.
@@ Appendix H Additional Experimental Results
In this section, we present complementary baseline results over all pretrained models that are not shown in the main text due to space limit. Table 1 is a summary of all pre-trained models we test on and their corresponding result tables for reference. Also, a complete table including all few-shot fine-tuning results for DWiSE-FT (including Fewshot 500 case omitted in the main text) are in Table LABEL:table:new_method_appendix.
@@
Split Methods ClinTox BBBP Bace HIV MUV Sider Tox21 ToxCast Avg Avg-F Avg-R Self-supervised Pre-training (MoleculeSTM) Random Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS Scaffold Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS Size Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS Supervised Pre-training (Graphium-Large) Scaffold Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS Size Full-FT LP Surgical-FT LP-FT WiSE-FT -SP Feature-map BSS
Split Methods Self-supervised Pre-training (MoleculeSTM) Supervised Pre-training (Graphium-Large) Esol Lipo Malaria Cep Avg-R Avg-R∗ Esol Lipo Malaria Cep Avg-R Avg-R∗ Random Full-FT 2 4 LP 8 8 Surgical-FT 6 1 LP-FT 3 2 WiSE-FT 5 3 -SP 4 7 Feature-map 7 5 BSS 1 6 Scaffold Full-FT 5 4.50 5 LP 8 6.25 7 Surgical-FT 6 2.75 3 LP-FT 4 3.75 4 WiSE-FT 2 4.75 6 -SP 3 7.50 8 Feature-map 7 3.50 2 BSS 1 3.00 1 Size Full-FT 2 2.75 2 LP 8 7.75 8 Surgical-FT 6 4.50 5 LP-FT 3 3.75 4 WiSE-FT 5 2.75 1 -SP 4 5.25 6 Feature-map 7 6.50 7 BSS 1 2.75 3
@@
Split Methods Self-supervised Pre-training (MoleculeSTM) Supervised Pre-training (Graphium-Large) Esol Lipo Malaria CEP Avg-R Avg-R∗ Esol Lipo Malaria CEP Avg-R Avg-R∗ Fewshot-50 Random Full-FT 6 7 LP 8 8 Surgical-FT 7 4 LP-FT 4 1 WiSE-FT 5 3 -SP 1 5 Feature-map 3 2 BSS 2 6 Scaffold Full-FT 6 4.00 7 LP 8 3.75 1 Surgical-FT 7 4.00 2 LP-FT 4 5.00 8 WiSE-FT 5 4.25 4 -SP 2 5.50 6 Feature-map 3 4.75 5 BSS 1 4.25 3 Size Full-FT 7 3.50 4 LP 8 7.50 8 Surgical-FT 6 4.50 1 LP-FT 4 5.75 7 WiSE-FT 5 4.50 5 -SP 1 4.25 6 Feature-map 3 3.25 3 BSS 2 2.50 2 Fewshot-100 Random Full-FT 6 5.00 6 LP 8 4.75 4 Surgical-FT 4 2.25 2 LP-FT 7 6.00 7 WiSE-FT 5 3.75 3 -SP 1 5.00 5 Feature-map 2 1.75 1 BSS 3 7.00 8 Scaffold Full-FT 6 6.75 7 LP 8 3.25 3 Surgical-FT 5 2.50 1 LP-FT 7 6.00 8 WiSE-FT 4 4.50 5 -SP 1 5.75 6 Feature-map 3 2.75 2 BSS 2 4.25 4 Size Full-FT 7 5.25 6 LP 8 7.75 8 Surgical-FT 6 2.50 1 LP-FT 5 6.50 7 WiSE-FT 4 4.00 5 -SP 1 2.75 3 Feature-map 2 3.75 4 BSS 3 3.00 2 Fewshot-500 Random Full-FT 6 5.00 5 LP 8 6.25 7 Surgical-FT 7 1.50 1 LP-FT 5 3.75 2 WiSE-FT 4 3.50 3 -SP 1 5.75 8 Feature-map 3 4.25 4 BSS 2 5.50 6 Scaffold Full-FT 6 5.50 6 LP 8 4.50 3 Surgical-FT 5 2.00 2 LP-FT 7 4.75 4 WiSE-FT 4 5.50 5 -SP 1 5.75 7 Feature-map 3 1.00 1 BSS 2 3.00 8 Size Full-FT 7 6.00 7 LP 8 8.00 8 Surgical-FT 4 3.75 4 LP-FT 5 3.25 3 WiSE-FT 6 4.75 5 -SP 1 1.75 1 Feature-map 2 5.50 6 BSS 3 2.25 2
@@
| Method | Few-Shot 50 (Scaffold Split) | Few-Shot 50 (Size Split) | ||||||||||
| BACE (AUC) | SIDER (AUC) | Avg AUC | ESOL (RMSE) | LIPO (RMSE) | Avg R | BACE (AUC) | SIDER (AUC) | Avg AUC | ESOL (RMSE) | LIPO (RMSE) | Avg R | |
| WiSE-FT | ||||||||||||
| L2-SP | ||||||||||||
| TOP | ||||||||||||
| DWiSE-FT | ||||||||||||
@@
| Scaffold Split | clintox | bbbp | bace | hiv | sider | esol | lipo | malaria | cep |
| fewshot 50 | |||||||||
| fewshot 500 | |||||||||
| non-fewshot |